# Solving General-Utility Markov Decision Processes in the Single-Trial Regime with Online Planning

### v1

@@ B.1 Proof of Theorem 1
=$\overset{(a)}{=}\mathbb{E}\left[f(\mathrm{d}^{\pi_{\text{M}}}_{\mathrm{s}_{0},\mathrm{a}_{0},\mathrm{s}_{1},\mathrm{a}_{1},\ldots}(s^{1}))\Bigg{|}\begin{subarray}{c}\mathrm{s}_{0}\sim p_{0},\mathrm{a}_{0}\sim\pi_{0}(\cdot|\mathrm{s}_{0}),\\ \mathrm{s}_{1}\sim P^{\mathrm{a}_{0}}(\cdot|s_{0}),\mathrm{a}_{1}\sim\pi_{1}(\cdot|\mathrm{s}_{1}),\\ \mathrm{s}_{2}\sim P^{\mathrm{a}_{1}}(\cdot|s_{1}),\mathrm{a}_{2}\sim\pi_{2}(\cdot|\mathrm{s}_{2}),\ldots\end{subarray}\right]$
=$\overset{(b)}{=}\mathbb{E}\left[\mathbb{E}\left[f(\mathrm{d}^{\pi_{\text{M}}}_{\mathrm{s}_{0},\mathrm{a}_{0},\mathrm{s}_{1},\mathrm{a}_{1},\ldots}(s^{1}))\Bigg{|}\begin{subarray}{c}\mathrm{a}_{3}\sim\pi_{3}(\cdot|s^{0}),\\ \mathrm{s}_{4}\sim P^{\mathrm{a}_{3}}(\cdot|s^{0}),\ldots\end{subarray}\right]\Bigg{|}\begin{subarray}{c}\mathrm{s}_{0}\sim p_{0},\mathrm{a}_{0}\sim\pi_{0}(\cdot|\mathrm{s}_{0}),\\ \mathrm{s}_{1}\sim P^{\mathrm{a}_{0}}(\cdot|s_{0}),\mathrm{a}_{1}\sim\pi_{1}(\cdot|\mathrm{s}_{1}),\\ \mathrm{s}_{2}\sim P^{\mathrm{a}_{1}}(\cdot|s_{1}),\mathrm{a}_{2}\sim\pi_{2}(\cdot|\mathrm{s}_{2})\end{subarray}\right]$
@@
=$\overset{(a)}{=}\frac{1}{2}\pi_{1}(a^{1}|s^{0})\left(f\left((1-\gamma)\left(1+\gamma^{2}+c\frac{\gamma^{4}}{1-\gamma^{2}}\right)\right)+f\left((1-\gamma)\left(\gamma^{2}+c\frac{\gamma^{4}}{1-\gamma^{2}}\right)\right)\right)$
@@
=$\overset{(a)}{=}\frac{1}{2}f\left(\frac{1-\gamma}{1-\gamma^{4}}\right)+\frac{1}{2}f\left(\frac{(1-\gamma)\gamma^{2}}{1-\gamma^{4}}\right)$
@@
As can be seen, the expression on the left-hand side of the inequality above corresponds to a weighted combination (with weights $\pi_{1}(a^{1}|s^{0})$ and $1-\pi_{1}(a^{1}|s^{0})$) of components (i) and (ii). By resorting to a software for symbolic/algebraic computation such as Mathematica [Wolfram Research, ] it can be shown that $(i)>\frac{1}{2}f\left(\frac{1-\gamma}{1-\gamma^{4}}\right)+\frac{1}{2}f\left(\frac{(1-\gamma)\gamma^{2}}{1-\gamma^{4}}\right)=F_{1}(\pi_{\text{NM}})$ and $(ii)>\frac{1}{2}f\left(\frac{1-\gamma}{1-\gamma^{4}}\right)+\frac{1}{2}f\left(\frac{(1-\gamma)\gamma^{2}}{1-\gamma^{4}}\right)=F_{1}(\pi_{\text{NM}})$ for any $c\in[0,1]$ and $\gamma\in(0,1)$. We provide the snippets of the Mathematica code we used below. This implies that the weighted combination satisfies the inequality above for any $\pi_{1}(a^{1}|s^{0})\in[0,1]$ and the conclusion follows.
@@ B.2 Proof of Proposition 1
[…]
@@
[…]
@@
[…]
@@
For every GUMDP $\mathcal{M}_{f}$ with $L$-Lipschitz $f$ and $H\in\mathbb{N}$, if $\pi^{*}=\operatorname*{arg\,min}_{\pi\in\Pi_{\textnormal{NM}}}F_{1,H}(\pi)$, then it holds that $\mathcal{R}(\pi^{*})\leq 4L\gamma^{H}.$
@@
[…]
@@
Proposition 1 (Regret decomposition).
@@
$\mathcal{R}(\pi)\leq\underbrace{F_{1,H}(\pi)-\min_{\pi_{H}\in\Pi_{\textnormal{NM}}}\left\{F_{1,H}(\pi_{H})\right\}}_{\text{$=\mathcal{R}_{H}(\pi)$}}+8L\gamma^{H},$ (10)
where $\mathcal{R}_{H}(\pi)$ is the regret of policy $\pi$ under the single-trial truncated objective with horizon $H$.
@@
$\mathcal{R}(\pi)$ $=\mathbb{E}\left[f(\mathbf{d}^{\pi})\right]-\min_{\pi^{\prime}\in\Pi_{\text{NM}}}\mathbb{E}\left[f(\mathbf{d}^{\pi^{\prime}})\right]$
@@
[…]
@@ B.3 The occupancy MDP: Value and action-value functions
For a given policy $\pi_{\text{O}}\in\Pi_{\text{S}}$, the interaction between the agent and the occupancy MDP gives rise to a random process ($(\{\mathrm{s}_{0},\mathbf{o}_{0}\},\mathrm{a}_{0},\{\mathrm{s}_{1},\mathbf{o}_{1}\},\mathrm{a}_{1},\ldots,\{\mathrm{s}_{H},\mathbf{o}_{H}\})$ such that:
@@
    [$\mathbb{P}\left[\{\mathrm{s}_{0},\mathbf{o}_{0}\}=\{s_{0},{\boldsymbol{o}}_{0}\}\right]=p_{0,\text{O}}(\{s_{0},{\boldsymbol{o}}_{0}\})$
@@
    [$\mathbb{P}\left[\{\mathrm{s}_{t+1},\mathbf{o}_{t+1}\}=\{s^{\prime},{\boldsymbol{o}}^{\prime}\}|\{\mathrm{s}_{0},\mathbf{o}_{0}\},\mathrm{a}_{0},\ldots,\{\mathrm{s}_{t},\mathbf{o}_{t}\},\mathrm{a}_{t}\right]=P_{\text{O}}^{\mathrm{a}_{t}}(\{\mathrm{s}_{t},\mathbf{o}_{t}\},\{\mathrm{s}^{\prime},\mathbf{o}^{\prime}\})$
@@
We let $(\Omega_{\text{O}},\mathcal{F}_{\text{O}},\mathbb{P}_{\pi_{\text{O}}}^{O})$ be the probability space over the sequence of random variables ($(\{\mathrm{s}_{0},\mathbf{o}_{0}\},\mathrm{a}_{0},\{\mathrm{s}_{1},\mathbf{o}_{1}\},\mathrm{a}_{1},\ldots,\{\mathrm{s}_{H},\mathbf{o}_{H}\})$ that satisfies conditions 1-3 above. We write specific trajectories as $\omega_{\text{O}}\in\Omega_{\text{O}}$, with $\omega_{\text{O}}=(\{s_{0},{\boldsymbol{o}}_{0}\},a_{0},\{s_{1},{\boldsymbol{o}}_{1}\},a_{1},\ldots,\{s_{H},{\boldsymbol{o}}_{H}\})$. We highlight that the probability of a given trajectory $\omega_{\text{O}}\in\Omega_{\text{O}}$ under stationary policy $\pi_{\text{O}}\in\Pi_{\text{S}}$ can be calculated as
$\mathbb{P}^{\text{O}}_{\pi_{\text{O}}}(\omega_{\text{O}})$ $=p_{0,\text{O}}(\{s_{0},{\boldsymbol{o}}_{0}\})\cdot\pi_{\text{O}}(a_{0}|\{s_{0},{\boldsymbol{o}}_{0}\})\cdot P_{\text{O}}^{a_{0}}(\{\mathrm{s}_{0},\mathbf{o}_{0}\},\{\mathrm{s}_{1},\mathbf{o}_{1}\})\ldots$
$\qquad\qquad\cdot\pi_{\text{O}}(a_{H-1}|\{s_{H-1},{\boldsymbol{o}}_{H-1}\})\cdot P_{\text{O}}^{a_{H-1}}(\{s_{H-1},{\boldsymbol{o}}_{H-1}\},\{s_{H},{\boldsymbol{o}}_{H}\}).$
To streamline our notation we introduce the mapping $\sigma:\mathcal{S}\times\mathcal{O}\times\mathcal{A}\rightarrow\mathcal{O}$ that describes the evolution of component ${\boldsymbol{o}}$ of the state, defined as
$\left[\sigma(s,{\boldsymbol{o}},a)\right]_{s^{\prime},a^{\prime}}=\begin{cases}o(s,a)+\gamma^{t}&\text{if $s^{\prime}=s,a^{\prime}=a$},\\ o(s,a)&\text{otherwise},\end{cases}$
where $[o]_{s^{\prime},a^{\prime}}$ denotes the value of entry $s^{\prime},a^{\prime}$ for vector ${\boldsymbol{o}}$, i.e., $[o]_{s^{\prime},a^{\prime}}=o(s^{\prime},a^{\prime})$.
@@
$V^{\pi_{\text{O}}}_{t}(\{s,{\boldsymbol{o}}\})$ $=\mathbb{E}_{\pi_{\text{O}}}\left[\sum_{t^{\prime}=t}^{H}c_{\text{O}}(\{\mathrm{s}_{t^{\prime}},\mathbf{o}_{t^{\prime}}\})\bigg{|}\{\mathrm{s}_{t},\mathbf{o}_{t}\}=\{s,{\boldsymbol{o}}\}\right]$ (11)
$=\mathbb{E}_{\pi_{\text{O}}}\left[c_{\text{O}}(\{\mathrm{s}_{H},\mathbf{o}_{H}\})\big{|}\{\mathrm{s}_{t},\mathbf{o}_{t}\}=\{s,{\boldsymbol{o}}\}\right],$ (12)
and the optimal value function, for any $t\in\{0,\ldots,H\}$, as $V^{*}_{t}(\{s,{\boldsymbol{o}}\})=\min_{\pi_{\text{O}}\in\Pi_{\text{S}}^{\text{D}}}V^{\pi_{\text{O}}}_{t}(\{s,{\boldsymbol{o}}\})$. The action-value function under $\mathcal{M}_{\text{O}}$, for any $t\in\{0,\ldots,H-1\}$, is defined as
$Q^{\pi_{\text{O}}}_{t}(\{s,{\boldsymbol{o}}\},a)$ $=\mathbb{E}_{\pi_{\text{O}}}\left[\sum_{t^{\prime}=t}^{H}c_{\text{O}}(\{\mathrm{s}_{t^{\prime}},\mathbf{o}_{t^{\prime}}\})\bigg{|}\{\mathrm{s}_{t},\mathbf{o}_{t}\}=\{s,{\boldsymbol{o}}\},\mathrm{a}_{t}=a\right]$
$=\mathbb{E}_{\pi_{\text{O}}}\left[c_{\text{O}}(\{\mathrm{s}_{H},\mathbf{o}_{H}\})\big{|}\{\mathrm{s}_{t},\mathbf{o}_{t}\}=\{s,{\boldsymbol{o}}\},\mathrm{a}_{t}=a\right],$
and the optimal action-value function, for any $t\in\{0,\ldots,H-1\}$, as $Q^{*}_{t}(\{s,{\boldsymbol{o}}\},a)=\min_{\pi_{\text{O}}\in\Pi_{\text{S}}^{\text{D}}}Q^{\pi_{\text{O}}}_{t}(\{s,{\boldsymbol{o}}\},a)$. We emphasize again that subscript $t$ can be dropped from $V^{\pi_{\text{O}}}_{t}(\{s,{\boldsymbol{o}}\})$, $V^{*}_{t}(\{s,{\boldsymbol{o}}\})$, $Q^{\pi_{\text{O}}}_{t}(\{s,{\boldsymbol{o}}\},a)$ and $Q^{*}_{t}(\{s,{\boldsymbol{o}}\},a)$ as it can be inferred from ${\boldsymbol{o}}$. Finally, we note that value functions, optimal value functions and optimal action-value functions satisfy the following set of Bellman equations:
$V^{\pi_{\text{O}}}_{t}(\{s,{\boldsymbol{o}}\})$ $=\sum_{a\in\mathcal{A}}\pi_{\text{O}}(a|\{s,{\boldsymbol{o}}\})\left(\sum_{s^{\prime}\in\mathcal{S}}P^{a}(s^{\prime}|s)V^{\pi_{\text{O}}}_{t+1}(\{s^{\prime},\sigma(s,{\boldsymbol{o}},a)\})\right),\quad\forall t\in\{0,\ldots,H-1\}$
$V^{*}_{t}(\{s,{\boldsymbol{o}}\})$ $=\min_{a\in\mathcal{A}}\left\{\sum_{s^{\prime}\in\mathcal{S}}P^{a}(s^{\prime}|s)V^{*}_{t+1}(\{s^{\prime},\sigma(s,{\boldsymbol{o}},a)\})\right\},\quad\forall t\in\{0,\ldots,H-1\}$
$Q^{*}_{t}(\{s,{\boldsymbol{o}}\},a)$ $=\sum_{s^{\prime}\in\mathcal{S}}P^{a}(s^{\prime}|s)V^{*}_{t+1}(\{s^{\prime},\sigma(s,{\boldsymbol{o}},a)\}),\quad\forall t\in\{0,\ldots,H-1\}$
$Q^{*}_{t}(\{s,{\boldsymbol{o}}\},a)$ $=\sum_{s^{\prime}\in\mathcal{S}}P^{a}(s^{\prime}|s)\min_{a^{\prime}\in\mathcal{A}}\left\{Q^{*}_{t+1}(\{s^{\prime},\sigma(s,{\boldsymbol{o}},a)\},a^{\prime})\right\},\quad\forall t\in\{0,\ldots,H-2\}.$

### v2

@@ B.1 Proof of Theorem 1
$\overset{(a)}{=}\mathbb{E}\left[f(\mathrm{d}^{\pi_{\text{M}}}_{\mathrm{s}_{0},\mathrm{a}_{0},\mathrm{s}_{1},\mathrm{a}_{1},\ldots}(s^{1}))\Bigg|\begin{subarray}{c}\mathrm{s}_{0}\sim p_{0},\mathrm{a}_{0}\sim\pi_{0}(\cdot|\mathrm{s}_{0}),\\ \mathrm{s}_{1}\sim P^{\mathrm{a}_{0}}(\cdot|s_{0}),\mathrm{a}_{1}\sim\pi_{1}(\cdot|\mathrm{s}_{1}),\\ \mathrm{s}_{2}\sim P^{\mathrm{a}_{1}}(\cdot|s_{1}),\mathrm{a}_{2}\sim\pi_{2}(\cdot|\mathrm{s}_{2}),\ldots\end{subarray}\right]$
$\overset{(b)}{=}\mathbb{E}\left[\mathbb{E}\left[f(\mathrm{d}^{\pi_{\text{M}}}_{\mathrm{s}_{0},\mathrm{a}_{0},\mathrm{s}_{1},\mathrm{a}_{1},\ldots}(s^{1}))\Bigg|\begin{subarray}{c}\mathrm{a}_{3}\sim\pi_{3}(\cdot|s^{0}),\\ \mathrm{s}_{4}\sim P^{\mathrm{a}_{3}}(\cdot|s^{0}),\ldots\end{subarray}\right]\Bigg|\begin{subarray}{c}\mathrm{s}_{0}\sim p_{0},\mathrm{a}_{0}\sim\pi_{0}(\cdot|\mathrm{s}_{0}),\\ \mathrm{s}_{1}\sim P^{\mathrm{a}_{0}}(\cdot|s_{0}),\mathrm{a}_{1}\sim\pi_{1}(\cdot|\mathrm{s}_{1}),\\ \mathrm{s}_{2}\sim P^{\mathrm{a}_{1}}(\cdot|s_{1}),\mathrm{a}_{2}\sim\pi_{2}(\cdot|\mathrm{s}_{2})\end{subarray}\right]$
@@
$\overset{(a)}{=}\frac{1}{2}\pi_{1}(a^{1}|s^{0})\left(f\left((1-\gamma)\left(1+\gamma^{2}+c\frac{\gamma^{4}}{1-\gamma^{2}}\right)\right)+f\left((1-\gamma)\left(\gamma^{2}+c\frac{\gamma^{4}}{1-\gamma^{2}}\right)\right)\right)$
@@
$\overset{(a)}{=}\frac{1}{2}f\left(\frac{1-\gamma}{1-\gamma^{4}}\right)+\frac{1}{2}f\left(\frac{(1-\gamma)\gamma^{2}}{1-\gamma^{4}}\right)$
@@
As can be seen, the expression on the left-hand side of the inequality above corresponds to a weighted combination (with weights $\pi_{1}(a^{1}|s^{0})$ and $1-\pi_{1}(a^{1}|s^{0})$) of components (i) and (ii). By resorting to a software for symbolic/algebraic computation such as Mathematica (Wolfram Research, ) it can be shown that $(i)>\frac{1}{2}f\left(\frac{1-\gamma}{1-\gamma^{4}}\right)+\frac{1}{2}f\left(\frac{(1-\gamma)\gamma^{2}}{1-\gamma^{4}}\right)=F_{1}(\pi_{\text{NM}})$ and $(ii)>\frac{1}{2}f\left(\frac{1-\gamma}{1-\gamma^{4}}\right)+\frac{1}{2}f\left(\frac{(1-\gamma)\gamma^{2}}{1-\gamma^{4}}\right)=F_{1}(\pi_{\text{NM}})$ for any $c\in[0,1]$ and $\gamma\in(0,1)$. We provide the snippets of the Mathematica code we used below. This implies that the weighted combination satisfies the inequality above for any $\pi_{1}(a^{1}|s^{0})\in[0,1]$ and the conclusion follows.
@@ B.2 Proof of Proposition 1
$\left|f(\mathbf{d}^{\pi}_{\omega})-f(\mathbf{d}^{\pi,H}_{\omega})\right|$ $\overset{\text{(a)}}{\leq}L\left\|\mathbf{d}^{\pi}_{\omega}-\mathbf{d}^{\pi,H}_{\omega}\right\|_{1}$
$\overset{\text{(b)}}{=}L\left\|(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}\mathbf{d}^{\pi}_{\omega,t}-\frac{(1-\gamma)}{1-\gamma^{H}}\sum_{t=0}^{H-1}\gamma^{t}\mathbf{d}^{\pi}_{\omega,t}\right\|_{1}$
@@
$\overset{\text{(c)}}{\leq}L\left(\frac{(1-\gamma)}{1-\gamma^{H}}\sum_{t=0}^{H-1}\gamma^{t}\left\|(1-\gamma^{H})\mathbf{d}^{\pi}_{\omega,t}-\mathbf{d}^{\pi}_{\omega,t}\right\|_{1}+(1-\gamma)\sum_{t=H}^{\infty}\gamma^{t}\|\mathbf{d}^{\pi}_{\omega,t}\|_{1}\right)$
@@
$\overset{\text{(a)}}{\leq}\mathbb{E}\left[\left|f(\mathbf{d}^{\pi})-f(\mathbf{d}^{\pi,H})\right|\right]$
@@
For every GUMDP $\mathcal{M}_{f}$ with $L$-Lipschitz $f$ and $H\in\mathbb{N}$, if $\pi^{*}=\operatorname*{arg\,min}_{\pi\in\Pi_{\textnormal{NM}}}F_{1,H}(\pi)$, then it holds that $\mathrm{OptGap}(\pi^{*})\leq 4L\gamma^{H}.$
@@
$F_{1,H}(\pi^{*})-2L\gamma^{H}=\min_{\pi}F_{1,H}(\pi)-2L\gamma^{H}\overset{\text{(a)}}{\leq}\min_{\pi}F_{1}(\pi)\overset{\text{(b)}}{\leq}F_{1}(\pi^{*}),$
@@
Proposition 1 (Optimality gap decomposition).
@@
$\mathrm{OptGap}(\pi)\leq\underbrace{F_{1,H}(\pi)-\min_{\pi_{H}\in\Pi_{\textnormal{NM}}}\left\{F_{1,H}(\pi_{H})\right\}}_{\text{$=\mathrm{OptGap}_{H}(\pi)$}}+8L\gamma^{H},$ (10)
where $\mathrm{OptGap}_{H}(\pi)$ is the optimality gap of policy $\pi$ under the single-trial truncated objective with horizon $H$.
@@
$\mathrm{OptGap}(\pi)$ $=\mathbb{E}\left[f(\mathbf{d}^{\pi})\right]-\min_{\pi^{\prime}\in\Pi_{\text{NM}}}\mathbb{E}\left[f(\mathbf{d}^{\pi^{\prime}})\right]$
@@
$\overset{\text{(a)}}{\leq}\left|\mathbb{E}\left[f(\mathbf{d}^{\pi})\right]-\mathbb{E}\left[f(\mathbf{d}^{\pi_{H}^{*}})\right]\right|+\left|\mathbb{E}\left[f(\mathbf{d}^{\pi_{H}^{*}})\right]-\min_{\pi^{\prime}\in\Pi_{\text{NM}}}\mathbb{E}\left[f(\mathbf{d}^{\pi^{\prime}})\right]\right|$
$\overset{\text{(b)}}{\leq}\left|\mathbb{E}\left[f(\mathbf{d}^{\pi})\right]-\mathbb{E}\left[f(\mathbf{d}^{\pi_{H}^{*}})\right]\right|+4L\gamma^{H}$
$\overset{\text{(c)}}{\leq}\left|\mathbb{E}\left[f(\mathbf{d}^{\pi})\right]-\mathbb{E}\left[f(\mathbf{d}^{\pi,H})\right]\right|+\left|\mathbb{E}\left[f(\mathbf{d}^{\pi,H})\right]-\mathbb{E}\left[f(\mathbf{d}^{\pi_{H}^{*}})\right]\right|+4L\gamma^{H}$
$\overset{\text{(d)}}{\leq}2L\gamma^{H}+\left|\mathbb{E}\left[f(\mathbf{d}^{\pi,H})\right]-\mathbb{E}\left[f(\mathbf{d}^{\pi_{H}^{*}})\right]\right|+4L\gamma^{H}$
$\overset{\text{(e)}}{\leq}\left|\mathbb{E}\left[f(\mathbf{d}^{\pi,H})\right]-\mathbb{E}\left[f(\mathbf{d}^{\pi_{H}^{*},H})\right]\right|+\left|\mathbb{E}\left[f(\mathbf{d}^{\pi_{H}^{*},H})\right]-\mathbb{E}\left[f(\mathbf{d}^{\pi_{H}^{*}})\right]\right|+6L\gamma^{H}$
$\overset{\text{(f)}}{\leq}\left|\mathbb{E}\left[f(\mathbf{d}^{\pi,H})\right]-\mathbb{E}\left[f(\mathbf{d}^{\pi_{H}^{*},H})\right]\right|+8L\gamma^{H}$
@@ B.3 The occupancy MDP: Value and action-value functions
For a given policy $\pi_{\text{O}}\in\Pi_{\text{S}}$, the interaction between the agent and the occupancy MDP gives rise to a random process $(\{\mathrm{s}_{0},\mathbf{o}_{0}\},\mathrm{a}_{0},\{\mathrm{s}_{1},\mathbf{o}_{1}\},\mathrm{a}_{1},\ldots,\{\mathrm{s}_{H},\mathbf{o}_{H}\})$ such that:
@@
    $\mathbb{P}\left[\{\mathrm{s}_{0},\mathbf{o}_{0}\}=\{s_{0},{\bm{o}}_{0}\}\right]=p_{0,\text{O}}(\{s_{0},{\bm{o}}_{0}\})$
@@
    $\mathbb{P}\left[\{\mathrm{s}_{t+1},\mathbf{o}_{t+1}\}=\{s^{\prime},{\bm{o}}^{\prime}\}|\{\mathrm{s}_{0},\mathbf{o}_{0}\},\mathrm{a}_{0},\ldots,\{\mathrm{s}_{t},\mathbf{o}_{t}\},\mathrm{a}_{t}\right]=P_{\text{O}}^{\mathrm{a}_{t}}(\{\mathrm{s}_{t},\mathbf{o}_{t}\},\{\mathrm{s}^{\prime},\mathbf{o}^{\prime}\})$
@@
We let $(\Omega_{\text{O}},\mathcal{F}_{\text{O}},\mathbb{P}_{\pi_{\text{O}}}^{O})$ be the probability space over the sequence of random variables $(\{\mathrm{s}_{0},\mathbf{o}_{0}\},\mathrm{a}_{0},\{\mathrm{s}_{1},\mathbf{o}_{1}\},\mathrm{a}_{1},\ldots,\{\mathrm{s}_{H},\mathbf{o}_{H}\})$ that satisfies conditions 1-3 above. We write specific trajectories as $\omega_{\text{O}}\in\Omega_{\text{O}}$, with $\omega_{\text{O}}=(\{s_{0},{\bm{o}}_{0}\},a_{0},\{s_{1},{\bm{o}}_{1}\},a_{1},\ldots,\{s_{H},{\bm{o}}_{H}\})$. We highlight that the probability of a given trajectory $\omega_{\text{O}}\in\Omega_{\text{O}}$ under stationary policy $\pi_{\text{O}}\in\Pi_{\text{S}}$ can be calculated as
$\mathbb{P}^{\text{O}}_{\pi_{\text{O}}}(\omega_{\text{O}})$ $=p_{0,\text{O}}(\{s_{0},{\bm{o}}_{0}\})\cdot\pi_{\text{O}}(a_{0}|\{s_{0},{\bm{o}}_{0}\})\cdot P_{\text{O}}^{a_{0}}(\{\mathrm{s}_{0},\mathbf{o}_{0}\},\{\mathrm{s}_{1},\mathbf{o}_{1}\})\ldots$
$\qquad\qquad\cdot\pi_{\text{O}}(a_{H-1}|\{s_{H-1},{\bm{o}}_{H-1}\})\cdot P_{\text{O}}^{a_{H-1}}(\{s_{H-1},{\bm{o}}_{H-1}\},\{s_{H},{\bm{o}}_{H}\}).$
To streamline our notation we introduce the mapping $\sigma:\mathcal{S}\times\mathcal{O}\times\mathcal{A}\rightarrow\mathcal{O}$ that describes the evolution of component $\bm{o}$} of the state, defined as
$\left[\sigma(s,{\bm{o}},a)\right]_{s^{\prime},a^{\prime}}=\begin{cases}o(s,a)+\gamma^{t}&\text{if $s^{\prime}=s,a^{\prime}=a$},\\ o(s,a)&\text{otherwise},\end{cases}$
where $o]_{s^{\prime},a^{\prime}}$ denotes the value of entry $s^{\prime},a^{\prime}$ for vector $\bm{o}$}, i.e., $o]_{s^{\prime},a^{\prime}}=o(s^{\prime},a^{\prime})$.
@@
$V^{\pi_{\text{O}}}_{t}(\{s,{\bm{o}}\})$ $=\mathbb{E}_{\pi_{\text{O}}}\left[\sum_{t^{\prime}=t}^{H}c_{\text{O}}(\{\mathrm{s}_{t^{\prime}},\mathbf{o}_{t^{\prime}}\})\bigg|\{\mathrm{s}_{t},\mathbf{o}_{t}\}=\{s,{\bm{o}}\}\right]$ (11)
$=\mathbb{E}_{\pi_{\text{O}}}\left[c_{\text{O}}(\{\mathrm{s}_{H},\mathbf{o}_{H}\})\big|\{\mathrm{s}_{t},\mathbf{o}_{t}\}=\{s,{\bm{o}}\}\right],$ (12)
and the optimal value function, for any $t\in\{0,\ldots,H\}$, as $V^{*}_{t}(\{s,{\bm{o}}\})=\min_{\pi_{\text{O}}\in\Pi_{\text{S}}^{\text{D}}}V^{\pi_{\text{O}}}_{t}(\{s,{\bm{o}}\})$. The action-value function under $\mathcal{M}_{\text{O}}$, for any $t\in\{0,\ldots,H-1\}$, is defined as
$Q^{\pi_{\text{O}}}_{t}(\{s,{\bm{o}}\},a)$ $=\mathbb{E}_{\pi_{\text{O}}}\left[\sum_{t^{\prime}=t}^{H}c_{\text{O}}(\{\mathrm{s}_{t^{\prime}},\mathbf{o}_{t^{\prime}}\})\bigg|\{\mathrm{s}_{t},\mathbf{o}_{t}\}=\{s,{\bm{o}}\},\mathrm{a}_{t}=a\right]$
$=\mathbb{E}_{\pi_{\text{O}}}\left[c_{\text{O}}(\{\mathrm{s}_{H},\mathbf{o}_{H}\})\big|\{\mathrm{s}_{t},\mathbf{o}_{t}\}=\{s,{\bm{o}}\},\mathrm{a}_{t}=a\right],$
and the optimal action-value function, for any $t\in\{0,\ldots,H-1\}$, as $Q^{*}_{t}(\{s,{\bm{o}}\},a)=\min_{\pi_{\text{O}}\in\Pi_{\text{S}}^{\text{D}}}Q^{\pi_{\text{O}}}_{t}(\{s,{\bm{o}}\},a)$. We emphasize again that subscript $t$ can be dropped from $V^{\pi_{\text{O}}}_{t}(\{s,{\bm{o}}\})$, $V^{*}_{t}(\{s,{\bm{o}}\})$, $Q^{\pi_{\text{O}}}_{t}(\{s,{\bm{o}}\},a)$ and $Q^{*}_{t}(\{s,{\bm{o}}\},a)$ as it can be inferred from $\bm{o}$}. Finally, we note that value functions, optimal value functions and optimal action-value functions satisfy the following set of Bellman equations:
$V^{\pi_{\text{O}}}_{t}(\{s,{\bm{o}}\})$ $=\sum_{a\in\mathcal{A}}\pi_{\text{O}}(a|\{s,{\bm{o}}\})\left(\sum_{s^{\prime}\in\mathcal{S}}P^{a}(s^{\prime}|s)V^{\pi_{\text{O}}}_{t+1}(\{s^{\prime},\sigma(s,{\bm{o}},a)\})\right),\quad\forall t\in\{0,\ldots,H-1\}$
$V^{*}_{t}(\{s,{\bm{o}}\})$ $=\min_{a\in\mathcal{A}}\left\{\sum_{s^{\prime}\in\mathcal{S}}P^{a}(s^{\prime}|s)V^{*}_{t+1}(\{s^{\prime},\sigma(s,{\bm{o}},a)\})\right\},\quad\forall t\in\{0,\ldots,H-1\}$
$Q^{*}_{t}(\{s,{\bm{o}}\},a)$ $=\sum_{s^{\prime}\in\mathcal{S}}P^{a}(s^{\prime}|s)V^{*}_{t+1}(\{s^{\prime},\sigma(s,{\bm{o}},a)\}),\quad\forall t\in\{0,\ldots,H-1\}$
$Q^{*}_{t}(\{s,{\bm{o}}\},a)$ $=\sum_{s^{\prime}\in\mathcal{S}}P^{a}(s^{\prime}|s)\min_{a^{\prime}\in\mathcal{A}}\left\{Q^{*}_{t+1}(\{s^{\prime},\sigma(s,{\bm{o}},a)\},a^{\prime})\right\},\quad\forall t\in\{0,\ldots,H-2\}.$